If the pth,qth and rth terms of a G.P are a,b,c resp.Pove that:-


(a^(q-r))(b^(r-p))(c^(p-q))=1

Grade:12

2 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Aparna,

given,

a  = xyp-1  ,  b = xyq-1 and c = xyr-1   here,  x is first term and y is common ratio

as, G2 = AH we ahve,

b2=ac

xyq-1

aq-r = xq-r * y(p-1)(q-r)

br-p  = xr-p * y(q-1)(r-p)

cp-q = xp-q * y(p-q)(r-1)

multiply all 3 of them, we get RHS,

x0 * y0 = 1

henc proved.

 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay

 

 

 

 

Aniket Patra
13 years ago

Just multiply the terms..and find the pth,qth and rth term by using the formula.(take x as the first term)..And then put the values in the expression..u will get (abc)^0 i.e. 1.